A researcher reports that the average salary of College Deans is more than P63,000. A sample of 35 College Deans has a mean salary of P65, 700. At , test the claim that the College Deans earn more than P63,000 a month. The standard deviation of the population is P5,250.


1
Expert's answer
2021-06-02T18:05:09-0400

actually exact question ;

A researcher reports that the average salary of College Deans is more than P63,000. A sample of 35 College Deans has a mean salary of P65, 700. At "\alpha =0.05" , test the claim that the College Deans earn more than P63,000 a month. The standard deviation of the population is P5,250.




solution:

given:

"\mu =P63000, \space n=35,\space \bar x =P65700 ,\space \sigma =P5250\\"

since we are given population standard deviation, we use z-test


step1:

state the hypothesis and identify the claim

"H_0 : \mu \leq P63000\\
H_1: \mu > P63000, claim(one-tailed \space right)"

stpe 2:

the level of significance is "\alpha =0.05"

step 3:

determine the critical value using the table

"z_{critical}=+1.645"

step 4:

compute the one sample z test value using the formula

"z_{computed}=\frac{\bar x- \mu}{\frac{\sigma }{\sqrt{n}}}\\
z_{computed}=\frac{65700- 63000}{\frac{5250}{\sqrt{35}}}\\
z_{computed}=\frac{2700}{\frac{5250}{\sqrt{35}}}\\
z_{computed}=\frac{2700}{5250} {\sqrt{35}}\\
z_{computed}=3.043"

step 5:

decision rule ,compare the computed and critical value of z

"3.043>1.645
\\
Reject\space H_0 \space and \space accept \space H_1"

step 6;

There is enough evidence to support the claim that the monthly salary of the college deans is more than P63000